# CryptOracle: A Modular Framework to Characterize FHE

### v2

@@ Abstract
Privacy-preserving machine learning has become an important long-term pursuit in this era of artificial intelligence (AI). Fully Homomorphic Encryption (FHE) is a uniquely promising solution, offering provable privacy and security guarantees. Unfortunately, computational cost is impeding its mass adoption. Modern solutions are up to six orders of magnitude slower than plaintext execution. Understanding and reducing this overhead is essential to the advancement of FHE, particularly as the underlying algorithms evolve rapidly. This paper presents a detailed characterization of OpenFHE, a comprehensive open-source library for FHE, with a particular focus on the CKKS scheme due to its significant potential for AI and machine learning applications. We introduce CryptOracle, a modular evaluation framework comprising (1) a benchmark suite, (2) a hardware profiler, and (3) a predictive performance model. The benchmark suite encompasses OpenFHE kernels at three abstraction levels: workloads, microbenchmarks, and primitives. The profiler is compatible with standard and user-specified security parameters. CryptOracle monitors application performance, captures microarchitectural events, and logs power and energy usage for AMD and Intel systems. These metrics are consumed by a modeling engine to estimate runtime and energy efficiency across different configuration scenarios, with error geomean of for runtime and for energy. CryptOracle is open source, fully modular, and serves as a shared platform to facilitate the collaborative advancements of applications, algorithms, software, and hardware in FHE. The CryptOracle code can be accessed at this Github repository.
@@ 1 Introduction
Modern computation is increasingly offloaded to the cloud, where shared infrastructure and limited transparency heighten the risk of data compromise. Contemporary security frameworks routinely encrypt data in storage and during transmission; however, they ultimately require plaintext access for computation. This window of exposure allows adversaries to exploit memory access patterns [40, 44], leverage microarchitectural side channels [18, 43], or trigger latent bugs in complex software stacks [12, 20, 36], especially under the proliferation of machine learning based attacks [54, 29].
Fully Homomorphic Encryption (FHE) offers a compelling alternative by enabling computation on encrypted data without decryption [26]. However, this guarantee comes with a much higher cost as compared to other secure-computation paradigms, not just because of the compute intensivity but also due to the evolving nature of workloads in the FHE space. Consequently, the systems community has pursued aggressive optimizations to overcome these challenges, catalyzing a vibrant ecosystem of architectural and algorithmic innovations, spanning general-purpose GPU implementations [23] to highly specialized hardware accelerators [51, 52, 35, 34, 3].
Problem Statement. Though existing profilers [45, 31, 25] provide valuable insights, a gap remains in automating rapid characterization across abstraction levels, hardware vendors, and evaluation metrics. We highlight challenges as below.
@@
Complicated optimization space. Performance in FHE is influenced by interdependent factors such as ciphertext size, multiplicative depth, noise growth, parallelism and memory layout. For example, ciphertexts can exceed 50 MB [35] and increase memory pressure; bootstrapping, taking up to 1.5s [51], is needed for noise management. Hence navigating this space involves tuning multiple architectural and algorithmic parameters, making design-space exploration less intuitive.
@@
The Solution. We introduce CryptOracle, a modular framework for systematic characterization of FHE applications, as shown in Figure 1. Built atop the community-maintained OpenFHE library [6], CryptOracle lowers the barrier to reproducible evaluation for theorists, system architects, and application developers alike. It comprises:
@@
We devise a fast performance-prediction model that estimates runtime and energy from primitive-level statistics.
We validate the model experimentally, distill architectural insights, and illustrate how CryptOracle accelerates collaborative research on FHE algorithms and systems.
@@ 2.1 FHE Framework
Introduced by Gentry in 2009[26], FHE allows computations to be performed directly on encrypted data, enabling privacy-preserving computation in untrusted environments. Following Gentry’s original construction, which was initially kept theoretical due to prohibitive computational overhead, subsequent advancements have led to the development of multiple FHE schemes that improve efficiency and functionality. Notable schemes include BGV[10], BFV[22], FHEW[19], TFHE[16], and CKKS[15].
The increasing interest in deploying FHE in real-world applications has catalyzed the development of several FHE libraries, including Microsoft SEAL[53], PALISADE[47], HEAAN[15], Lattigo[1], HELayers[4] and OpenFHE[6], each supporting a certain number of different schemes. Among these, OpenFHE is one of the most comprehensive ones. First, OpenFHE supports a wide range of FHE scheme under a unified API. Second, it offers compatibility with the Homomorphic Encryption Standard[37], ensuring cryptographic soundness and reproducibility. Third, OpenFHE allows fine-grained customization of encryption parameters, enabling performance-security tradeoff studies across varied hardware backends. Finally, it features built-in support for both leveled and bootstrapped CKKS, facilitating benchmarking across arbitrary depths. OpenFHE supports a rich set of homomorphic operations, including ciphertext-plaintext and ciphertext-ciphertext addition and multiplication, as well as advanced primitives such as vector rotation, scalar encoding, conjugation, bootstrapping.
@@ 2.2 The CKKS Scheme
CKKS leverages the Number Theoretic Transform (NTT) to accelerate polynomial multiplication: once mapped to the NTT domain, the operation reduces to a point-wise product, reducing the computational complexity from to . Each homomorphic multiplication amplifies the accumulated noise and simultaneously scales the plaintext. To curb this noise growth, the ciphertext is rescaled by dividing it by its last modulus prime [14]: , where . After several such multiply–rescale rounds, the remaining modulus eventually becomes too small to accommodate further multiplications, which caps the achievable circuit depth. Because CKKS does not natively support non-linear functions such as ReLU, computations rely on polynomial approximations, which typically demand a substantial multiplicative depth[38]. To permit computations of arbitrary depth, CKKS employs bootstrapping[26]. This procedure homomorphically evaluates the decryption followed by the re-encryption of a ciphertext, thereby resetting its noise budget and restoring both its modulus and available multiplicative depth. Consequently, encrypted data can undergo an unlimited sequence of operations while still producing correct results. Bootstrapping is indispensable, yet it remains computationally intensive and consumes several levels of the ciphertext modulus. State-of-the-art implementations accelerate bootstrapping through homomorphic FFTs and polynomial approximations[13], making it the focus of ongoing research in both hardware and algorithmic optimizations.
@@ 2.3 Benchmarking and Characterization
Benchmarking and workload characterization are essential for understanding application behavior and for guiding both system- and compiler-level optimizations. Over time, diverse methodologies have been devised to capture the performance characteristics of both general-purpose and domain-specific workloads. General-purpose benchmark suites such as SPEC CPU2017 [11], Graph500 [27], and MLPerf [42] have driven advances in evaluating compute-intensive, irregular-memory-acess, and deep-learning applications. In parallel, specialized characterization frameworks have been proposed to target particular system behaviors. Bircher[9], for instance, leveraged hardware performance counters for comprehensive system-level power estimation, whereas Crape[17] developed a statistically rigorous methodology for profiling dynamic Python workloads. Liu[41] further extended this direction by proposing a data-efficient framework tailored to domain-specific architectures. Despite their breadth, these studies concentrate on conventional, plaintext applications and therefore overlook the unique challenges of FHE, whose encrypted computation, intricate algebraic operations, and unique performance bottlenecks demand specialized profiling strategies.
@@ 2.4 Motivation of CryptOracle
CryptOracle benchmark suite is built on OpenFHE, targeting CKKS across three abstraction levels: workload, microbenchmark, and primitive. The hardware profiler captures CPU behaviors at the primitive level and reports a diverse set of hardware metrics. Using the runtime and energy measured for primitives, the performance model predicts the corresponding values for workloads and microbenchmarks.
Recent advances in FHE software and hardware acceleration have moved FHE closer to practical deployment [21, 46, 3, 52]. Even with this progress, the community still lacks a unified, systematic framework for performance benchmarking.
@@
Where suites such as SPEC or MLPerf evaluate general-purpose programs, FHE workloads intertwine cryptographic kernels with high-dimensional algebraic routines and operations that map poorly to commodity hardware. Consequently, CryptOracle profiles execution at the primitive level. This approach, common in high-performance computing and accelerator design, leverages the consistent and predictable behavior of primitives to guide optimizations across all layers. Profiling with primitive granularity exposes system bottlenecks in fine detail, enabling targeted optimization [32, 23].
@@
The high computational cost of FHE has motivated extensive work on custom hardware accelerators[51, 52, 35, 34, 3]. These efforts usually appear alongside new applications and algorithms. Because implementing each idea on actual hardware is complex and time-consuming, assessing the benefits only after deployment is impractical. A faster approach is to estimate performance and efficiency on general-purpose CPUs through analytical or empirical models. Similar modeling has become standard practice for predicting the runtime of AI workloads at various scales[48, 30, 39]. Following this methodology, CryptOracle estimates the performance and energy of FHE applications from primitive-level profiling, enabling accurate yet rapid exploration of algorithmic and application innovations.
@@ 3.1 Overview
| Thread count: maximum allowable | ||
| OpenMP threads | ||
@@
| Performance | Target FHE abstraction: microbenchmark | |
@@
| Output | Callstack graph, profiling results, | |
| prediction results | ||
@@ 3.3.2 Event Profiling
3.3.3 Statistics Denoising
After profiling, the raw statistics recorded during the setup phase are subtracted from those collected during the steady-state execution phase for each timing, energy, and perf counter, isolating the program’s region of interest (ROI). For low-latency primitives, these values are averaged across all internal calls to ensure a sufficiently large sample. Next, average power consumption is calculated. When the instructions and cpu-cycles perf events are enabled, the application’s IPC is calculated. All metrics are stored for subsequent data analysis and performance model construction.
| Category | Name | Description |
| Primitive | EvalAdd | Adds two ciphertexts |
| EvalAdd (Plaintext) | Adds a ciphertext and a plaintext | |
| EvalSub | Difference of two ciphertexts | |
| EvalSub (Scalar) | Difference of a scalar from every element of a ciphertext | |
| EvalMult | Multiplication of two ciphertexts | |
| EvalMultNoRelin | Multiplication of two ciphertexts without relinearization | |
| EvalMult (Plaintext) | Multiplication of a ciphertext and a plaintext | |
| EvalMult (Scalar) | Multiplication of every element of a ciphertext by a scalar | |
| EvalSquare | Multiplication of a ciphertext with itself | |
| EvalRotate | Rotates an encrypted vector by a specified index | |
| EvalFastRotate | Rotates using precomputed automorphisms | |
| EvalBootstrap | Refreshes a ciphertext to a predefined level count | |
| EvalChebyshevFunction | Evaluates a smooth function over [a,b] using Chebyshev polynomials | |
| EvalChebyshevSeries | Evaluates a Chebyshev polynomial series over [a,b] with specified coefficients | |
| Microbenchmark | Matrix Multiplication | Multiplies two square ciphertext matrices |
| Logistic Function | Chebyshev approximation of the logistic function evaluated on an encrypted vector | |
| Sign Evaluation Function | Chebyshev approximation of the sign(x) function on an encrypted vector | |
| Workload | CIFAR-10 Image Classification | Uses a pre-trained CIFAR-10 model to classify encrypted images |
| ResNet20 Image Classification | Uses a pre-trained ResNet-20 model to classify encrypted images | |
| Logistic Regression Training | Performs ten iterations of logistic regression training | |
| Chi Square Test Evaluation | Implements the chi-square test for genome-wide association studies (GWAS) |
| Primitive | Matrix Multiplication | Logistic Function | Sign Evaluation | CIFAR-10 | ResNet-20 | Logistic Regression | Chi-Squared Test |
| EvalAdd | 178 | 70 | 519 | 47 | 6845 | 39 | 598 |
| EvalAdd (Plaintext) | 0 | 0 | 0 | 2 | 24 | 0 | 0 |
| EvalSub | 0 | 85 | 310 | 3 | 0 | 19 | 7 |
| EvalSub (Scalar) | 0 | 31 | 256 | 4 | 0 | 0 | 0 |
| EvalMult | 16 | 116 | 568 | 7 | 219 | 9 | 0 |
| EvalMultNoRelin | 0 | 0 | 0 | 0 | 0 | 0 | 207 |
| EvalMult (Plaintext) | 32 | 0 | 0 | 33 | 6475 | 40 | 4 |
| EvalMult (Scalar) | 0 | 67 | 61 | 0 | 0 | 0 | 3 |
| EvalSquare | 0 | 8 | 0 | 0 | 0 | 0 | 0 |
| EvalRotate | 193 | 0 | 0 | 33 | 1218 | 20 | 0 |
| EvalFastRotate | 0 | 0 | 0 | 0 | 152 | 0 | 0 |
| EvalBootstrap | 0 | 0 | 0 | 0 | 21 | 9 | 0 |
| EvalChebyshevFunction | 0 | 0 | 0 | 0 | 12 | 0 | 0 |
| EvalChebyshevSeries | 0 | 1 | 1 | 0 | 0 | 0 | 0 |
@@ 3.3.4 Callstack Recording
During a third pass of the hardware profiler, users can record the application’s call stack. Capturing this information reveals which FHE operations dominate execution time. The resulting stack trace is automatically rendered as a FlameGraph [28]. In a FlameGraph, the width of each box indicates the proportion of total runtime spent in the corresponding function, while the vertical axis represents call-stack depth. Figure 4 presents the logistic function, showing that roughly 90% of its runtime is consumed by the homomorphic multiply operation. This visualization enables developers to identify performance bottlenecks at early stages.
@@ 3.4 Performance Model
Packaged with the hardware profiler is a lightweight performance model that rapidly evaluates application runtime and other metrics for security parameters and thread counts that have not yet been tested. We use application to denote both microbenchmarks and full workloads. The model receives two inputs: the primitive profiling results and the counts of every homomorphic operation in the target application (for a list of common homomorphic operations, see Table 2). It then linearly extrapolates the application runtime (e.g. ResNet-20 inference) by summing, for each primitive, the product of its count and its measured runtime at the chosen security parameters and thread count. Because FHE abstractions exhibit long runtimes, this linear approach provides projections with minimal overhead. This model could then be easily used to evaluate the impact of algorithmic changes to primitive operations or scaling an application to a higher security level. To assess the model’s fidelity, CryptOracle includes a routine that validates the estimated runtime and reports the estimation error.

### v3

@@ Abstract
Privacy-preserving machine learning has become an important long-term pursuit in this era of artificial intelligence (AI). Fully Homomorphic Encryption (FHE) is a uniquely promising solution, offering provable privacy and security guarantees. Unfortunately, computational cost is impeding its mass adoption. Modern solutions are up to six orders of magnitude slower than plaintext execution. Understanding and reducing this overhead is essential to the advancement of FHE, particularly as the underlying algorithms evolve rapidly. This paper presents a detailed characterization of OpenFHE, a comprehensive open-source library for FHE, with a particular focus on the CKKS scheme due to its significant potential for AI and machine learning applications. We introduce CryptOracle, a modular evaluation framework comprising (1) a benchmark suite, (2) a hardware profiler, and (3) a predictive performance model. The benchmark suite encompasses OpenFHE kernels at three abstraction levels: workloads, microbenchmarks, and primitives. The profiler is compatible with standard and user-specified security parameters. CryptOracle monitors application performance, captures microarchitectural events, and logs power and energy usage for AMD and Intel systems. These metrics are consumed by a modeling engine to estimate runtime and energy efficiency across different configuration scenarios, with prediction error ranging from for runtime and for energy (geomean). CryptOracle is open source, fully modular, and serves as a shared platform to facilitate the collaborative advancements of applications, algorithms, software, and hardware in FHE. The CryptOracle code can be accessed at this anonymous Github repository.
@@ 1 Introduction
Modern computation is increasingly offloaded to the cloud, where shared infrastructure and limited transparency heighten the risk of data compromise. Contemporary security frameworks routinely encrypt data in storage and during transmission; however, they ultimately require plaintext access for computation. This window of exposure allows adversaries to exploit memory access patterns [memoryTraceOblivious, camouflage], leverage microarchitectural side channels [svf, osvikcache2005], or trigger latent bugs in complex software stacks [foreshadow, heartbleed, spectre], especially under the proliferation of machine learning based attacks [MembershipAttack, GANAttack].
Fully Homomorphic Encryption (FHE) offers a compelling alternative by enabling computation on encrypted data without decryption [gentry2009fully]. However, this guarantee comes with a much higher cost as compared to other secure-computation paradigms, not just because of the computational intensity but also due to the evolving nature of workloads in the FHE space. Consequently, the systems community has pursued aggressive optimizations to overcome these challenges, catalyzing a vibrant ecosystem of architectural and algorithmic innovations, spanning general-purpose GPU implementations [fan2023tensorfhe] to highly specialized hardware accelerators [samardzic2021f1, samardzic2022craterlake, kim2022bts, kim2022ark, agrawal2023fab].
Problem Statement. Though existing profilers [pal2023fully, jiang2022fhebench, hebench] provide valuable insights, a gap remains in automating rapid characterization across abstraction levels, hardware vendors, and evaluation metrics. We highlight challenges as below.
@@
Complicated optimization space. Performance in FHE is influenced by interdependent factors such as ciphertext size, multiplicative depth, noise growth, parallelism and memory layout. For example, ciphertexts can exceed 50 MB [kim2022bts] and increase memory pressure; bootstrapping, taking up to 1.5s [samardzic2021f1], is needed for noise management. Hence navigating this space involves tuning multiple architectural and algorithmic parameters, making design-space exploration less intuitive.
@@
The Solution. We introduce CryptOracle, a modular framework for systematic characterization of FHE applications, as shown in Figure 1. Built atop the community-maintained OpenFHE library [OpenFHE], CryptOracle lowers the barrier to reproducible evaluation for theorists, system architects, and application developers alike. It comprises:
@@
We devise a fast prediction model that estimates performance from primitive-level statistics.
We validate the model experimentally, distill architectural insights, and illustrate how CryptOracle accelerates research on FHE algorithms and systems.
@@ 2.1 FHE Framework
Introduced by Gentry in 2009[gentry2009fully], FHE allows computations to be performed directly on encrypted data, enabling privacy-preserving computation in untrusted environments. Following Gentry’s original construction, which was initially kept theoretical due to prohibitive computational overhead, subsequent advancements have led to the development of multiple FHE schemes that improve efficiency and functionality. Notable schemes include BGV[brakerski2014leveled], BFV[fan2012somewhat], FHEW[ducas2015fhew], TFHE[chillotti2020tfhe], and CKKS[cheon2017homomorphic].
The increasing interest in deploying FHE in real-world applications has catalyzed the development of several FHE libraries, including Microsoft SEAL[sealcrypto], PALISADE[polyakov2017palisade], HEAAN[cheon2017homomorphic], Lattigo[lattigo], HELayers[aharoni2020helayers] and OpenFHE[OpenFHE], each supporting a certain number of different schemes. Among these, OpenFHE is one of the most comprehensive ones. First, OpenFHE supports a wide range of FHE scheme under a unified API. Second, it offers compatibility with the Homomorphic Encryption Standard[lauter2022protecting], ensuring cryptographic soundness and reproducibility. Third, OpenFHE allows fine-grained customization of encryption parameters, enabling performance-security tradeoff studies across varied hardware backends. Finally, it features built-in support for both leveled and bootstrapped CKKS, facilitating benchmarking across arbitrary depths. OpenFHE supports a rich set of homomorphic operations, including ciphertext-plaintext and ciphertext-ciphertext addition and multiplication, as well as advanced primitives such as vector rotation, scalar encoding, conjugation, bootstrapping.
@@ 2.2 The CKKS Scheme
CKKS leverages the Number Theoretic Transform (NTT) to accelerate polynomial multiplication: once mapped to the NTT domain, the operation reduces to a point-wise product, reducing the computational complexity from to . Each homomorphic multiplication amplifies the accumulated noise and simultaneously scales the plaintext. To curb this noise growth, the ciphertext is rescaled by dividing it by its last modulus prime [cheon2018full]: , where . After several multiply–rescale rounds, the remaining modulus eventually becomes too small to accommodate further multiplications, capping the achievable circuit depth. Because CKKS does not natively support nonlinear functions (e.g. ReLU), computations rely on polynomial approximations, which typically demand a substantial multiplicative depth[lee2023precise]. To permit computations of arbitrary depth, CKKS employs bootstrapping[gentry2009fully]. This procedure homomorphically evaluates the decryption followed by the re-encryption of a ciphertext, thereby resetting its noise budget and restoring both its modulus and available multiplicative depth. Consequently, encrypted data can undergo an unlimited sequence of operations while still producing correct results. Bootstrapping is indispensable, yet it remains computationally intensive and consumes several levels of the ciphertext modulus. State-of-the-art implementations accelerate bootstrapping through FFTs and polynomial approximations[cheon2018bootstrapping], making it the focus of ongoing research in hardware and algorithmic optimizations.
@@ 2.3 Benchmarking and Characterization
Benchmarking and workload characterization are essential for understanding application behavior and for guiding both system- and compiler-level optimizations. Over time, diverse methodologies have been devised to capture the performance characteristics of both general-purpose and domain-specific workloads. General-purpose benchmark suites such as SPEC CPU2017 [spec_cpu2017], Graph500 [graph500], and MLPerf [mattson2020mlperftrainingbenchmark] have driven advances in evaluating compute-intensive, irregular-memory-acess, and deep-learning applications. In parallel, specialized characterization frameworks have been proposed to target particular system behaviors. Bircher[bircher2011complete], for instance, leveraged hardware performance counters for comprehensive system-level power estimation, whereas Crape[crape2020rigorous] developed a statistically rigorous methodology for profiling dynamic Python workloads. Liu[liu2020deffe] further extended this direction by proposing a data-efficient framework tailored to domain-specific architectures. Despite their breadth, these studies concentrate on conventional, plaintext applications and therefore overlook the unique challenges of FHE, whose encrypted computation, intricate algebraic operations, and unique performance bottlenecks demand specialized profiling strategies.
@@ 2.4 Motivation of CryptOracle
CryptOracle is built on OpenFHE, targeting CKKS across three abstraction levels: workload, microbenchmark, and primitive. The hardware profiler captures CPU behaviors at the primitive level and reports a diverse set of hardware metrics. Using the runtime and energy measured for primitives, the performance model predicts the corresponding values for workloads and microbenchmarks.
Recent advances in FHE software and hardware acceleration have moved FHE closer to practical deployment [ebel2025orion, park2025fhendi, agrawal2023fab, samardzic2022craterlake]. Even with this progress, the community still lacks a unified, systematic framework for benchmarking.
@@
Where suites such as SPEC or MLPerf evaluate general-purpose programs, FHE workloads intertwine cryptographic kernels with high-dimensional algebraic routines and operations that map poorly to commodity hardware. Consequently, CryptOracle profiles execution at the primitive level. This approach, common in HPC and accelerator design, leverages the consistent and predictable behavior of primitives to guide optimizations across all layers. Profiling with primitive granularity exposes system bottlenecks in fine detail, enabling targeted optimization [jung2021over, fan2023tensorfhe].
@@
The high computational cost of FHE has motivated extensive work on custom hardware accelerators[samardzic2021f1, samardzic2022craterlake, kim2022bts, kim2022ark, agrawal2023fab]. These efforts usually appear alongside new applications and algorithms. Because implementing each idea on actual hardware is complex and time-consuming, assessing the benefits only after deployment is impractical. A faster approach is to estimate performance and efficiency on general-purpose CPUs through analytical or empirical models. Similar modeling has become standard practice for predicting the runtime of AI workloads at various scales[qi17paleo, calculon_paper, 10.1145/3669940.3707265]. Following this methodology, CryptOracle estimates the performance and energy of FHE applications from primitive-level profiling, enabling accurate yet rapid exploration of algorithmic and application innovations.
@@ 3.1 Overview
| Thread count: max allowable OMP threads | ||
@@
| Performance | Target abstraction: microbenchmark | |
@@
| Output | Callstack graph, profiling/prediction results | |
@@ 3.2 Benchmark Suite
| Category | Application | Sec. Std. | N | B | L |
| Primitive | OpenFHE Primitives | none | 10 | ||
| Micro- | Matrix Multiplication | none | 10 | ||
| Logistic Function | none | 10 | |||
| benchmark | Sign Evaluation | none | 10 | ||
| Workload | CIFAR-10 Inference | none | 5 | ||
| Chi Square Test | none | 1 | 3 | ||
| ResNet20 Inference | none | 10 | |||
| Logreg Training | 128-bit | 14 |
@@ 3.3.2 Event Profiling
3.3.3 Statistics Denoising
After profiling, the raw statistics recorded during the setup phase are subtracted from those collected during the steady-state execution phase for each timing, energy, and perf counter, isolating the program’s region of interest (ROI). For low-latency primitives, these values are averaged across all internal calls to ensure a sufficiently large sample. Next, average power consumption is calculated.
All metrics are stored for subsequent data analysis and performance model construction.
@@ 3.3.4 Callstack Recording
During a third pass of the hardware profiler, users can record the application’s call stack. Capturing this information reveals which FHE operations dominate execution time. The resulting stack trace is automatically rendered as a FlameGraph [gregg_flamegraph_2017]. In a FlameGraph, the width of each box indicates the proportion of total runtime spent in the corresponding function, while the vertical axis represents call-stack depth. Figure 4 presents the logistic function, showing that roughly 90% of its runtime is consumed by the homomorphic multiply operation. This visualization enables developers to identify performance bottlenecks at early stages.
@@ 3.4 Performance Model
Packaged with the hardware profiler is a lightweight performance model that rapidly evaluates application runtime and other metrics for security parameters and thread counts that have not yet been tested. We use application to denote both microbenchmarks and full workloads. The model receives two inputs: the primitive profiling results and the counts of every homomorphic operation in the target application. It then linearly extrapolates the application runtime (e.g. ResNet-20 inference) by summing, for each primitive, the product of its count and its measured runtime at the chosen security parameters and thread count. Because FHE abstractions exhibit long runtimes, this linear approach provides projections with minimal overhead. This model could then be easily used to evaluate the impact of algorithmic changes to primitive operations or scaling an application to a higher security level. To assess the model’s fidelity, CryptOracle includes a routine that validates the estimated runtime and reports the estimation error.